# InfoTok: Adaptive Discrete Video Tokenizer via Information-Theoretic Compression

### v1

@@ 1 Introduction
With the growing success of vision foundation models (Agarwal et al., 2025; Zhang et al., 2025), there has been an increasing interest to model the real world with visual observations, e.g., videos. Representing video as discrete tokens offers clear advantages by aligning seamlessly with LLMs and unifying diverse vision tasks, yet discrete video tokenization remains highly challenging, especially for complex visual content. The main difficulty lies in developing efficient and scalable representations for long sequences, which can easily expand to millions of tokens and become intractable for the transformer architecture (Esser et al., 2021; Wang et al., 2024; Agarwal et al., 2025).
A discrete tokenizer typically consists of an encoder, a quantizer, and a decoder. Given a video with a certain duration and resolution, the encoder compresses the original video into a sequence of embeddings, which are subsequently quantized into discrete tokens. During decoding, the decoder reconstructs the video from these token sequences. The training objective aims to minimize the discrepancy between the original and reconstructed videos—an objective achievable if token sequences are allowed to be arbitrarily long (e.g., directly preserving all RGB pixel values). Hence, an effective tokenizer inherently acts as a compressor, encoding visual signals into compact tokens while preserving essential information for downstream tasks. Unlike images (Van Den Oord et al., 2017; Mentzer et al., 2023b; Tian et al., 2024; Yu et al., 2024), video signals pose unique tokenization challenges due to their high dimensionality, temporal redundancy, and considerable variability in information content (Yan et al., 2024; Agarwal et al., 2025). Therefore, developing an efficient, scalable, and principled video tokenizer is crucial for advancing unified multi-modal models.
As is well established in Shannon information theory (Ash, 2012), compared with content-agnostic compression that compresses all data identically, content-dependent compression, which determines the compression ratio of a data sample according to its property, is significantly more efficient. However, regardless of architectures and quantization methods, most existing tokenizers use an identical compression rate for all possible videos (Yan et al., 2021; Esser et al., 2021; Yu et al., 2023a). Consequently, a fixed amount of tokens could be redundant for simple videos and insufficient for representing complex videos, making downstream video-understanding or generation tasks inefficient or intractable. While recent attempts on flexible tokenization (Yan et al., 2024) use a heuristic training method that enables users to tokenize a video with varying lengths, we prove in this paper that their training method is biased, and their inference via trial-and-error length selection is inefficient. As a result, we seek to answer the following principled and challenging problem:
@@
To answer the question, this paper draws inspiration from Shannon’s Source Coding Theorem (Shannon, 1948) and proves that existing tokenizers, either with a fixed compression rate or data-agnostic adaptive rates, are biased in the sense that the expected token length is significantly larger than the optimal length in order to achieve identical reconstruction quality. In contrast, tokenizing each video in a difficulty-aware manner, i.e., which is closely tied to the frequency of contents based on information theory, can guarantee a near-optimal compression rate in theory. Altogether, our theoretical results suggest a principled way for adaptive tokenization.
@@ 2.1 Discrete Video Tokenization
where $\mathbb{D}$} is the video distribution to sample $\mathbf{x}$} from, and $q_{\phi}({\mathbf{z}}|{\mathbf{x}})$ and $p_{\theta}({\mathbf{x}}|{\mathbf{z}})$ are the posteriors for compression and reconstruction parameterized by the encoder ${\mathcal{E}}_{\phi}$ and decoder ${\mathcal{D}}_{\theta}$, respectively. Under the standard assumption of VAE (Kingma et al., 2019), eq.˜1 can be converted into a MSE loss on the pixel space. As such, tokenizers are essentially compressors that compress $\mathbf{x}$} into a dense discrete sequence $\mathbf{z}$}.
@@ 2.2 A Unified Framework for Adaptive Tokenization
In this section, we analyze the inherent limitations of fixed-compression discrete tokenizers. We show that such a constraint is sub-optimal due to the varying spatio-temporal complexity across videos. Specifically, different videos contain diverse spatial content (e.g., scenes, objects) and exhibit varying temporal dynamics (e.g., speed, magnitude of motion), resulting in unequal informational demands. From the perspective of information theory, frequent contents are naturally easier to represent than infrequent contents, leading to a natural result of varying representation lengths. To better illustrate this sub-optimality, consider an idealized scenario where the tokenizer $\mathcal{T}$ can perfectly reconstruct any input video $\mathbf{x}$} with non-zero probability $p({\mathbf{x}})>0$. In what follows, we restate the Shannon Source Coding Theorem (Shannon et al., 1959) from Information Theory to ground our argument.
Theorem 2.1 (Shannon Source Coding Theorem (restated) (Shannon et al., 1959)).
@@ 2.3 Sub-optimality of Existing Methods
Specifying $N_{\mathbf{x}}$ for each video is challenging because adaptiveness is not only for flexibility but also for principled compression. As special cases of our framework, several existing tokenizers for images (Duggal et al., 2024) and videos (Duggal et al., 2024) propose to train with a heuristic and data-agnostic router, where $N_{\mathbf{x}}$ is sampled uniformly from $\{1,\cdots,N\}$. During inference, the sequence length is selected based on the quality. However, uniform selection does not take into account the different information quantities possessed by different videos. Furthermore, the uncertainty in random selection brings additional overhead at inference time: To determine the optimal length, one needs to perform a search over possible lengths. To rigorously prove why this is biased, we simplify their inference stage and assume that an oracle $r^{*}(N_{\mathbf{x}}|{\mathbf{x}})$ can directly return the minimal sequence length conditioned on the loss eq.˜2 being minimized. The complete proof is deferred to appendix˜A.
@@ 3.1 Information-theoretic Adaptive Tokenization
where $\beta$ is the average compression factor that measures how much information a token should represent, $\delta(\cdot)$ denotes the delta distribution, and the term $\mathbb{E}[\mathrm{ELBO}({\mathbf{x}})]$ is for normalization. In addition to the theoretical merit that is discussed below, our choice directly approximates the optimal sequence length through eq.˜4 and thus does not require extensive probing over different lengths during both training and inference as Yan et al. (2024). In practice, eq.˜4 can be computed efficiently: we first encode $\mathbf{x}$} and decode back to $\hat{\mathbf{x}}$ without using the adaptive compressor, so as to compute the reconstruction error, which is the negative ELBO when adding the additional KL term. We then reuse the continuous embedding $\mathbf{h}$} and then compute ${\mathbf{h}}^{\prime}={\mathcal{M}}_{\psi}({\mathbf{h}},N_{\mathbf{x}})$, which is henceforth quantized into a token sequence $\mathbf{z}$} with length $N_{\mathbf{x}}$. As a result, we only have one additional decoder pass in order to generate the compressed token sequence.
@@ 3.2 Adaptive Compressor
Architecture. As a general framework, InfoTok can be seamlessly integrated on top of established tokenizer architectures by reusing their encoder ${\mathcal{E}}_{\phi}$ and decoder ${\mathcal{D}}_{\theta}$. Empirically, we implement InfoTok based on Cosmos Discrete Video Tokenizer (Agarwal et al., 2025), a 3D-CNN-based fixed-length tokenizer used to initialize ${\mathcal{E}}_{\phi}$ and ${\mathcal{D}}_{\theta}$. To account for the extra flexibility introduced in the adaptive compressor and decompressor, we additionally initialize the compressor ${\mathcal{M}}_{\psi}$ with a stack of multiple Transformer layers to effectively compress $\mathbf{h}$} into $\mathbf{h}$$^{\prime}$. Similarly, the decompressor contains multiple Transformer layers for mapping the condensed representation back to the full-length sequence.
@@ 4.1 Experimental Settings
Dataset. To systematically evaluate different tokenization, we consider two video reconstruction datasets, namely TokenBench proposed in Agarwal et al. (2025), and DAVIS in Caelles et al. (2019), each containing videos with various resolutions and numbers of frames. Since ElasticTok only takes in square 256px video with $H=W=256$, we only take the 256px partition of both datasets and crop each video to a square shape randomly. All results below are reported on our processed datasets for fair comparisons, although this will make baseline results not directly comparable with their paper. Notably, InfoTok can generalize to video with different resolutions, as presented in Appendix. C.
Evaluations. We evaluate the quality of reconstructed videos by four metrics: the Peak Signal to Noise Ratio (PSNR) (Hore & Ziou, 2010), the Structural Similarity (SSIM) (Hore & Ziou, 2010), the Learned Perceptual Image Patch Similarity (LPIPS) metric (Zhang et al., 2018), and the Fréchet Video Distance (Unterthiner et al., 2019). To evaluate the compression, we adopt the bits-per-pixel (BPP) value that computes how many digital bits are used to represent one pixel (with all color channels) in the whole video. For the sake of presentation, we present BPP with a unit of $\frac{1}{16}$, which is represented as $\text{BPP}_{16}$ (bits-per-16-pixels). For instance, if the number of tokens is $c\cdot THW$ for a video with resolution $H\times W$ and length $T$, and the codebook size is $C$, then $\text{BPP}_{16}=16c\cdot\log(C)$.
InfoTok Details. We use Cosmos Tokenizer (Agarwal et al., 2025) as the fixed-length tokenizer and use it to compute the ELBO value of inputs. Given a video of shape $T\times H\times W$, it uses $N_{\text{max}}=\frac{T}{4}\times\frac{H}{8}\times\frac{W}{8}=\frac{1}{256}THW$ tokens to tokenize the video. We refer our methods using the router $r_{\beta}(N_{\mathbf{x}}|{\mathbf{x}})$ as InfoTok and the one using $r^{\text{flex}}_{\beta}(N_{\mathbf{x}}|{\mathbf{x}})$ as InfoTok-Flex. For router $r_{\beta}(N_{\mathbf{x}}|{\mathbf{x}})$, $\mathbb{E}[\mathrm{ELBO}({\mathbf{x}})]$ is computed from the exponential moving average of historical training samples. For router $r_{\beta}^{\text{flex}}(N_{\mathbf{x}}|{\mathbf{x}})$, we consider ${\mathcal{B}}=\{0.25N_{\text{max}},0.5N_{\text{max}},0.75N_{\text{max}},N_{\text{max}}\}$ to assemble tokenizers with different average compression rates. During inference, we specify an average $\text{BPP}_{16}$, and $\beta$ can be computed as $N_{\text{max}}\cdot(\text{BPP}_{16}-\frac{1}{16})$, where $\frac{1}{16}$ is the cost of binary mask. The adaptive compressor is an eight-layer transformer with a block-causal attention matrix that maintains the causality of the Cosmos Tokenizer. For quantization, we use the Finite Scalar Quantizer (Mentzer et al., 2023b).
@@
Baselines. First, for fixed-compression tokenizers, we evaluate Open‐MAGVIT2 (codebook size $C=2^{18}$) (Yu et al., 2023a), OmniTokenizer ($C=2^{13}$) (Wang et al., 2024), and the Cosmos Discrete Video Tokenizer ($C=2^{16}$) (Agarwal et al., 2025). Their resulting $\text{BPP}_{16}$ values are reported in Table 1. For adaptive tokenization, we compare against ElasticTok ($2^{16}$) (Yan et al., 2024), which employs right-to-left random masking on token sequence during training and binary search to select the minimal number of tokens to meet a specified reconstruction‐loss threshold, with $\text{BPP}_{16}$ ranging from $0.25$ to $4$. Since ElasticTok’s adaptiveness is governed by a loss threshold rather than an average compression rate, we align our methods with their settings.
@@ 4.3 Ablation Study
Table 3: Ablation on adaptive compressors. All results are in average $\text{BPP}_{16}=0.56$.
@@
Ablation on adaptive compressor. To better understand our adaptive compressor that leverages ELBO for compression, we compare it with two additional strategies that are intuitive: simple right-to-left masking used by ElasticTok (R2L), and a spatially dispersed variant that masks a token every four tokens (Jump), in Table 3. Our compressor that masks tokens by ELBO values yields better reconstruction quality. These findings demonstrate that the compressor design matters in adaptive tokenization.
Table 4: Ablation on different variants of adaptive mechanisms across architectures. All results are obtained with average $\text{BPP}_{16}=0.56$ on TokenBench.
@@
Ablation on the adaptive mechanism. Unifying the router and adaptive compression, we further compare the whole adaptive mechanism of InfoTok to ElasticTok. Specifically, we evaluate both the adaptive mechanism of ElasticTok and InfoTok, to the Cosmos architecture, and a pure ViT architecture. As shown in Table 4, InfoTok consistently outperforms ElasticTok across both architectures, yielding substantial improvements in both PSNR and FVD. These results confirm the effectiveness and generalization ability of InfoTok.
@@ 5 Related Works
Discrete Tokenization has been extensively explored to facilitate generative AI. Typical VQ-VAE (Van Den Oord et al., 2017) and VQGAN (Esser et al., 2021) extend the Variational Autoencoder (VAE) framework (Kingma et al., 2013) by introducing a learnable codebook to quantize 2D encoded features into discrete tokens, which are then decoded back to reconstruct the original image. To streamline the quantization process, lookup-free tokenization methods such as FSQ (Mentzer et al., 2023b) and LFQ (Yu et al., 2023c) have been proposed. In terms of representation, TiTok (Yu et al., 2024) employs a Transformer-based architecture to learn 1D tokens. Building upon this, FlowMo (Sargent et al., 2025) and SelfTok (Pan et al., 2025) utilize diffusion decoders conditioned on 1D discrete tokens to mitigate discretization errors further. Along the line of video tokenization, there are several works, including OminiTokenizer (Wang et al., 2024), Open-MAGVIT2 (Luo et al., 2024), and Cosmos (Agarwal et al., 2025). Unlike images, video data is intrinsically more complex and sparse, making existing tokenization approaches inefficient.
Adaptive Representation has been studied over the past decade, with early works like Nested Dropout (Rippel et al., 2014) and Matryoshka Representation Learning (Kusupati et al., 2022) introducing methods to learn representations at multiple granularities. Recent approaches have applied similar methodologies to tokenization. CAT (Shen et al., 2025) focuses on continuous tokens, using vision-language models to assess image content complexity heuristically to determine token usage. ALIT (Duggal et al., 2024) recurrently distills 2D tokens into 1D latent tokens, enabling flexible tokenization. One-D-Piece (Miwa et al., 2025) and ElasticTok (Yan et al., 2024) employ random nested dropout strategies for tokens, targeting images and videos, respectively. FlexTok (Bachmann et al., 2025) further incorporates a diffusion decoder to enhance reconstruction quality under low token counts. However, these works focus on images and rely on heuristic methods for adaptive tokenization such as random masking. As such, their objective functions are biased by definition.

### v2

@@ 1 Introduction
With the growing success of vision foundation models (agarwal2025cosmos; zhang2025videollama), there has been an increasing interest to model the real world with visual observations, e.g., videos. Representing video as discrete tokens offers clear advantages by aligning seamlessly with LLMs and unifying diverse vision tasks, yet discrete video tokenization remains highly challenging, especially for complex visual content. The main difficulty lies in developing efficient and scalable representations for long sequences, which can easily expand to millions of tokens and become intractable for the transformer architecture (esser2021taming; wang2024omnitokenizer; agarwal2025cosmos).
A discrete tokenizer typically consists of an encoder, a quantizer, and a decoder. Given a video with a certain duration and resolution, the encoder compresses the original video into a sequence of embeddings, which are subsequently quantized into discrete tokens. During decoding, the decoder reconstructs the video from these token sequences. The training objective aims to minimize the discrepancy between the original and reconstructed videos—an objective achievable if token sequences are allowed to be arbitrarily long (e.g., directly preserving all RGB pixel values). Hence, an effective tokenizer inherently acts as a compressor, encoding visual signals into compact tokens while preserving essential information for downstream tasks. Unlike images (van2017vqvae; mentzer2023fsq; tian2024var; yu2024titok), video signals pose unique tokenization challenges due to their high dimensionality, temporal redundancy, and considerable variability in information content (yan2024elastictok; agarwal2025cosmos). Therefore, developing an efficient, scalable, and principled video tokenizer is crucial for advancing unified multi-modal models.
As is well established in Shannon information theory (ash2012information), compared with content-agnostic compression that compresses all data identically, content-dependent compression, which determines the compression ratio of a data sample according to its property, is significantly more efficient. However, regardless of architectures and quantization methods, most existing tokenizers use an identical compression rate for all possible videos (yan2021videogpt; esser2021taming; yu2023magvit). Consequently, a fixed amount of tokens could be redundant for simple videos and insufficient for representing complex videos, making downstream video-understanding or generation tasks inefficient or intractable. While recent attempts on flexible tokenization (yan2024elastictok) use a heuristic training method that enables users to tokenize a video with varying lengths, we prove in this paper that their training method is biased, and their inference via trial-and-error length selection is inefficient. As a result, we seek to answer the following principled and challenging problem:
@@
To answer the question, this paper draws inspiration from Shannon’s Source Coding Theorem (shannon1948mathematical) and proves that existing tokenizers, either with a fixed compression rate or data-agnostic adaptive rates, are biased in the sense that the expected token length is significantly larger than the optimal length in order to achieve identical reconstruction quality. In contrast, tokenizing each video in a difficulty-aware manner, i.e., which is closely tied to the frequency of contents based on information theory, can guarantee a near-optimal compression rate in theory. Altogether, our theoretical results suggest a principled way for adaptive tokenization.
@@ 2.1 Discrete Video Tokenization
where $\mathbb{D}$} is the video distribution to sample $\mathbf{x}$} from, and $q_{\phi}({\mathbf{z}}|{\mathbf{x}})$ and $p_{\theta}({\mathbf{x}}|{\mathbf{z}})$ are the posteriors for compression and reconstruction parameterized by the encoder ${\mathcal{E}}_{\phi}$ and decoder ${\mathcal{D}}_{\theta}$, respectively. Under the standard assumption of VAE (kingma2019introduction), eq.˜1 can be converted into a MSE loss on the pixel space. As such, tokenizers are essentially compressors that compress $\mathbf{x}$} into a dense discrete sequence $\mathbf{z}$}.
@@ 2.2 A Unified Framework for Adaptive Tokenization
In this section, we analyze the inherent limitations of fixed-compression discrete tokenizers. We show that such a constraint is sub-optimal due to the varying spatio-temporal complexity across videos. Specifically, different videos contain diverse spatial content (e.g., scenes, objects) and exhibit varying temporal dynamics (e.g., speed, magnitude of motion), resulting in unequal informational demands. From the perspective of information theory, frequent contents are naturally easier to represent than infrequent contents, leading to a natural result of varying representation lengths. To better illustrate this sub-optimality, consider an idealized scenario where the tokenizer $\mathcal{T}$ can perfectly reconstruct any input video $\mathbf{x}$} with non-zero probability $p({\mathbf{x}})>0$. In what follows, we restate the Shannon Source Coding Theorem (shannon1959coding) from Information Theory to ground our argument.
Theorem 2.1 (Shannon Source Coding Theorem (restated) (shannon1959coding)).
@@ 2.3 Sub-optimality of Existing Methods
Specifying $N_{\mathbf{x}}$ for each video is challenging because adaptiveness is not only for flexibility but also for principled compression. As special cases of our framework, several existing tokenizers for images (duggal2024alit) and videos (duggal2024alit) propose to train with a heuristic and data-agnostic router, where $N_{\mathbf{x}}$ is sampled uniformly from $\{1,\cdots,N\}$. During inference, the sequence length is selected based on the quality. However, uniform selection does not take into account the different information quantities possessed by different videos. Furthermore, the uncertainty in random selection brings additional overhead at inference time: To determine the optimal length, one needs to perform a search over possible lengths. To rigorously prove why this is biased, we simplify their inference stage and assume that an oracle $r^{*}(N_{\mathbf{x}}|{\mathbf{x}})$ can directly return the minimal sequence length conditioned on the loss eq.˜2 being minimized. The complete proof is deferred to appendix˜A.
@@ 3.1 Information-theoretic Adaptive Tokenization
where $\beta$ is the average compression factor that measures how much information a token should represent, $\delta(\cdot)$ denotes the delta distribution, and the term $\mathbb{E}[\mathrm{ELBO}({\mathbf{x}})]$ is for normalization. In addition to the theoretical merit that is discussed below, our choice directly approximates the optimal sequence length through eq.˜4 and thus does not require extensive probing over different lengths during both training and inference as yan2024elastictok. In practice, eq.˜4 can be computed efficiently: we first encode $\mathbf{x}$} and decode back to $\hat{\mathbf{x}}$ without using the adaptive compressor, so as to compute the reconstruction error, which is the negative ELBO when adding the additional KL term. We then reuse the continuous embedding $\mathbf{h}$} and then compute ${\mathbf{h}}^{\prime}={\mathcal{M}}_{\psi}({\mathbf{h}},N_{\mathbf{x}})$, which is henceforth quantized into a token sequence $\mathbf{z}$} with length $N_{\mathbf{x}}$. As a result, we only have one additional decoder pass in order to generate the compressed token sequence.
@@ 3.2 Adaptive Compressor
Architecture. As a general framework, InfoTok can be seamlessly integrated on top of established tokenizer architectures by reusing their encoder ${\mathcal{E}}_{\phi}$ and decoder ${\mathcal{D}}_{\theta}$. Empirically, we implement InfoTok based on Cosmos Discrete Video Tokenizer (agarwal2025cosmos), a 3D-CNN-based fixed-length tokenizer used to initialize ${\mathcal{E}}_{\phi}$ and ${\mathcal{D}}_{\theta}$. To account for the extra flexibility introduced in the adaptive compressor and decompressor, we additionally initialize the compressor ${\mathcal{M}}_{\psi}$ with a stack of multiple Transformer layers to effectively compress $\mathbf{h}$} into $\mathbf{h}$$^{\prime}$. Similarly, the decompressor contains multiple Transformer layers for mapping the condensed representation back to the full-length sequence.
@@ 4.1 Experimental Settings
Dataset. To systematically evaluate different tokenization, we consider two video reconstruction datasets, namely TokenBench proposed in agarwal2025cosmos, and DAVIS in sergi2019davis, each containing videos with various resolutions and numbers of frames. Since ElasticTok only takes in square 256px video with $H=W=256$, we only take the 256px partition of both datasets and crop each video to a square shape randomly. All results below are reported on our processed datasets for fair comparisons, although this will make baseline results not directly comparable with their paper. Notably, InfoTok can generalize to video with different resolutions, as presented in Appendix. C.
Evaluations. We evaluate the quality of reconstructed videos by four metrics: the Peak Signal to Noise Ratio (PSNR) (hore2010psnrssim), the Structural Similarity (SSIM) (hore2010psnrssim), the Learned Perceptual Image Patch Similarity (LPIPS) metric (zhang2018lpips), and the Fréchet Video Distance (unterthiner2019fvd). To evaluate the compression, we adopt the bits-per-pixel (BPP) value that computes how many digital bits are used to represent one pixel (with all color channels) in the whole video. For the sake of presentation, we present BPP with a unit of $\frac{1}{16}$, which is represented as $\text{BPP}_{16}$ (bits-per-16-pixels). For instance, if the number of tokens is $c\cdot THW$ for a video with resolution $H\times W$ and length $T$, and the codebook size is $C$, then $\text{BPP}_{16}=16c\cdot\log(C)$.
InfoTok Details. We use Cosmos Tokenizer (agarwal2025cosmos) as the fixed-length tokenizer and use it to compute the ELBO value of inputs. Given a video of shape $T\times H\times W$, it uses $N_{\text{max}}=\frac{T}{4}\times\frac{H}{8}\times\frac{W}{8}=\frac{1}{256}THW$ tokens to tokenize the video. We refer our methods using the router $r_{\beta}(N_{\mathbf{x}}|{\mathbf{x}})$ as InfoTok and the one using $r^{\text{flex}}_{\beta}(N_{\mathbf{x}}|{\mathbf{x}})$ as InfoTok-Flex. For router $r_{\beta}(N_{\mathbf{x}}|{\mathbf{x}})$, $\mathbb{E}[\mathrm{ELBO}({\mathbf{x}})]$ is computed from the exponential moving average of historical training samples. For router $r_{\beta}^{\text{flex}}(N_{\mathbf{x}}|{\mathbf{x}})$, we consider ${\mathcal{B}}=\{0.25N_{\text{max}},0.5N_{\text{max}},0.75N_{\text{max}},N_{\text{max}}\}$ to assemble tokenizers with different average compression rates. During inference, we specify an average $\text{BPP}_{16}$, and $\beta$ can be computed as $N_{\text{max}}\cdot(\text{BPP}_{16}-\frac{1}{16})$, where $\frac{1}{16}$ is the cost of binary mask. The adaptive compressor is an eight-layer transformer with a block-causal attention matrix that maintains the causality of the Cosmos Tokenizer. For quantization, we use the Finite Scalar Quantizer (mentzer2023fsq).
@@
Baselines. First, for fixed-compression tokenizers, we evaluate Open‐MAGVIT2 (codebook size $C=2^{18}$) (yu2023magvit), OmniTokenizer ($C=2^{13}$) (wang2024omnitokenizer), and the Cosmos Discrete Video Tokenizer ($C=2^{16}$) (agarwal2025cosmos). Their resulting $\text{BPP}_{16}$ values are reported in Table 1. For adaptive tokenization, we compare against ElasticTok ($2^{16}$) (yan2024elastictok), which employs right-to-left random masking on token sequence during training and binary search to select the minimal number of tokens to meet a specified reconstruction‐loss threshold, with $\text{BPP}_{16}$ ranging from $0.25$ to $4$. Since ElasticTok’s adaptiveness is governed by a loss threshold rather than an average compression rate, we align our methods with their settings.
@@ 4.3 Ablation Study
Table 3: Ablation results on TokenBench with an average $\text{BPP}_{16}=0.56$. (Left) Ablation on adaptive compressors. (Right) Ablation on different variants of adaptive mechanisms across architectures.
@@
Ablation on adaptive compressor. To better understand the ELBO-based compression, we compare it with two additional intuitive strategies: simple right-to-left masking used by ElasticTok (R2L), and a spatially dispersed variant that masks a token every four tokens (Jump), in Table 3 (Left). Our compressor that masks tokens by ELBO values yields better reconstruction quality. These findings demonstrate that the compressor design matters in adaptive tokenization.
Ablation on the adaptive mechanism. Unifying the router and adaptive compression, we further compare the whole adaptive mechanism of InfoTok to ElasticTok. Specifically, we evaluate both the adaptive mechanism of ElasticTok and InfoTok, to the Cosmos architecture, and a pure ViT architecture. As shown in Table 3 (Right), InfoTok consistently outperforms ElasticTok across both architectures, yielding substantial improvements in both PSNR and FVD. These results confirm the effectiveness and generalization ability of InfoTok.
@@ 5 Related Works
Discrete Tokenization has been extensively explored to facilitate generative AI. Typical VQ-VAE (van2017vqvae) and VQGAN (esser2021taming) extend the Variational Autoencoder (VAE) framework (kingma2013vae) by introducing a learnable codebook to quantize 2D encoded features into discrete tokens, which are then decoded back to reconstruct the original image. To streamline the quantization process, lookup-free tokenization methods such as FSQ (mentzer2023fsq) and LFQ (yu2023lfq) have been proposed. In terms of representation, TiTok (yu2024titok) employs a Transformer-based architecture to learn 1D tokens. Building upon this, FlowMo (sargent2025flowmo) and SelfTok (pan2025selftok) utilize diffusion decoders conditioned on 1D discrete tokens to mitigate discretization errors further. Along the line of video tokenization, there are several works, including OminiTokenizer (wang2024omnitokenizer), Open-MAGVIT2 (luo2024open), and Cosmos (agarwal2025cosmos). Unlike images, video data is intrinsically more complex and sparse, making existing tokenization approaches inefficient.
Adaptive Representation has been studied over the past decade, with early works like Nested Dropout (rippel2014learning) and Matryoshka Representation Learning (kusupati2022matryoshka) introducing methods to learn representations at multiple granularities. Recent approaches have applied similar methodologies to tokenization. CAT (shen2025cat) focuses on continuous tokens, using vision-language models to assess image content complexity heuristically to determine token usage. ALIT (duggal2024alit) recurrently distills 2D tokens into 1D latent tokens, enabling flexible tokenization. One-D-Piece (miwa2025one) and ElasticTok (yan2024elastictok) employ random nested dropout strategies for tokens, targeting images and videos, respectively. FlexTok (bachmann2025flextok) further incorporates a diffusion decoder to enhance reconstruction quality under low token counts. However, these works focus on images and rely on heuristic methods for adaptive tokenization such as random masking. As such, their objective functions are biased by definition.